Complete the following for aqueous solutions of Aluminum nitrate. Show Complete Solution.

  1. Mass of Solute= 2.544g, Molarity= 1.688M, Volume of Solution=____
  2. Mass of Solute= _____, Molarity= 0.729M, Volume of Solution= 894mL
1
Expert's answer
2022-03-04T04:57:02-0500

Aluminium nitrate= 213

1) moles= 2.544/213= 0.012

Volume= (0.012/1.688)× 1000= 7.1ml

2)

(894/1000)× 0.729= 0.6517moles

Mass= 213× 0.6517= 138.8g